# Coincidences of Division Fields of an elliptic curve defined over a number field

### v2

@@ Introduction
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{4}$
@@
Let $E/F$ be an elliptic curve with an $(m,n)$-coincidence. Let $p$ be a prime such that $p\mid n$ and $p\nmid m$. Let $\mathfrak{p}$ an ideal above $p$ such that the index of ramification of $F/\mathbb{Q}$ at $\mathfrak{p}$ is prime to $\varphi(p^{k})$. Then we are in one of the following cases:
@@ 1. Notations and preliminaries
    For a positive integer $m$, we set $\mathrm{GL}_{2}(m):=\mathrm{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$ and $\mathrm{SL}_{2}(m):=\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z})$.
    For finite extensions $K\subseteq L\subseteq M$ such that $M/L$ is Galois, and a prime ideal $\mathfrak{p}$ of $L$, we denote by $e_{\mathfrak{p}}(M/K)$ the index of ramification of $M/K$ at $\mathfrak{p}$.
@@
First of all, we prove Proposition 1.2 given in the introduction after some preliminaries. Let $E/F$ be an elliptic curve. If $E/F$ has CM, we suppose that $F$ contains the CM field. From Serre’s open image theorem [Ser72, Section 4.4, Theorem 3’], there exists a finite set $S$ of rational primes such that
@@
Let $m\geq 2$, $p$ be a prime such that $p>q$ for all primes $q\mid m$ and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\zeta_{p^{\infty}})$. Let $E/F$ be an elliptic curve such that $F(\zeta_{p^{k}})\subseteq F(E[m])$ with $k>r$. Then, $k=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case either $r=0$ and $k\leq 2$, or $r=k-1$.
@@ 2.2. Ramification and entanglement
The extension $F(\zeta_{p^{k}})/F$ is Galois and so the index of ramification above $\mathfrak{p}$ only depends on $\mathfrak{p}$. We have
@@
additive red. at $\mathfrak{p}$, $p>3$ $v_{p}(t)=0$
@@
multiplicative red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$
@@
add. red. at $\mathfrak{p}$, $p>3$ $v_{p}(t)=0$ Proposition 2.3
@@
mult. red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$ Proposition 2.2
@@
By Lemma 2.5, we obtain the table below, in which we present the necessary condition on the ramification of $F/\mathbb{Q}$ to obtain the index of ramification as in previous theorem.
@@
With the notation of Theorem 2.6 and Remark 2.7, if $F(\zeta_{p^{k}})\subseteq F(E[m])$, then we must have $s\mid t$. Therefore, the tables give restrictions on having $F(\zeta_{p^{k}})\subseteq F(E[m])$. For example, if we have this inclusion and $E/F$ has good reduction at $\mathfrak{p}$, then $\varphi(p^{k})$ must divide the index of ramification of $F/\mathbb{Q}$ at $\mathfrak{p}$. In the following corollary, we consider the case of $F/\mathbb{Q}$ unramified above $p$.
@@ 3.1. Construction of vertical coincidences
$1\to\operatorname{\mathrm{Gal}}(F(E[mn]/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn]/F)\to\operatorname{\mathrm{Gal}}(F(E[m]/F)\to 1$
@@
    $1\to\operatorname{\mathrm{Gal}}(F(E[mn]/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn]/F)\to\operatorname{\mathrm{Gal}}(F(E[m]/F)\to 1$
@@
$1\to(\mathbb{Z}/2\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8]/\mathbb{Q})\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[4]/\mathbb{Q})\to 1$
@@
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{r}$
for some $r\leq 4$. Indeed, the Galois group $\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))$ is isomorphic, for $n=2$, to a subgroup of
@@
$1\to(\mathbb{Z}/p\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F)\to\operatorname{\mathrm{Gal}}(F(E[p^{k}]/F)\to 1$
@@ 3.2. Trivial intersection with the cyclotomic field
If $E/F$ has an $(m,mn)$-coincidence, then we must have $F(\zeta_{mn})\subseteq F(E[m])$. But this does not implies in general that $F(\zeta_{mn})=F(\zeta_{m})$, as we will see in Remark 3.24. Even more, unless $m$ is odd and $n=2$, this last never happens if $F=\mathbb{Q}$, and yet some coincidences occurs, like $(2,4)$ and $(2,6)$-coincidence, see [DLR23, Examples 1.2 and 1.3]. As in Remark 3.24, it is due to the non-surjectivity of $\mathrm{SL}_{2}(mn)\cap\rho_{E,mn}(G_{F})\to\mathrm{SL}_{2}(m)\cap\rho_{E,m}(G_{F})$.
@@ 3.3. Index of images
Proposition 3.16.
Let $E/F$ be an elliptic curve without CM, with a $(p^{k},p^{k+1})$-coincidence. Then $p^{4}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$.
Proof.
With the introduced notation, we consider $M=\mathrm{GL}_{2}(\mathbb{Z}_{p})$ and $G=\rho_{E,p^{\infty}}(G_{F})$. The index $i_{k+1}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ by Remark 3.5 and so $i_{k+1}/i_{k}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$. But having a $(p^{k},p^{k+1})$-coincidence is equivalent to have $G_{k}\simeq G_{k+1}$. Then, the proposition follows from Equation 4. ∎
Remark 3.17.
In [Zyw22, Theorem 1.3], Zywina gives a set of possible index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{\mathbb{Q}})]$ which is generic for elliptic curves $E$ defined over $\mathbb{Q}$. In [Zyw24, Theorem 1.2], he gives upper bounds for the index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ for elliptic curves $E/F$, depending on the number field $F$.
@@
    If $G=\mathrm{GL}_{2}(\mathbb{Z}_{p})$, then $\left(\frac{i_{k+1}}{i_{k}}\right)$ is constant equal to $1$. The converse is false: the elliptic curve with LMFDB label 11.a2 has non maximal Galois representation at $5$ and the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ attached to $5$ is constant, equal to $1$.
  3. (3)
    For $E/\mathbb{Q}$ without CM, the $2$-adic image is the inverse image in $\mathrm{GL}_{2}(\mathbb{Z}_{2})$ of $\rho_{E,32}(G_{\mathbb{Q}})$, from [RZB15, Corollary 1.3]. Therefore, the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ stabilizes at $1$ from $k$ at most $5$.
@@
Here some examples over $\mathbb{Q}$, computed from [LMF24], illustrate different possibilities for the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$:
@@
In the first case, we have seen in Remark 3.23 that $F(\zeta_{p^{k+1}})$ is equal to $F(\zeta_{4})$ if $p=2$ and $F(\zeta_{p})$ otherwise. The examples of vertical coincidence we have for elliptic curves over $\mathbb{Q}$ fits, obviously, in the second case. Indeed, the elliptic curve with LMFDB label 40.a4 has a $(2,4)$-coincidence with
@@ 3.4. Split liftable subgroups
Now we will present groups which are not split liftable, which give us obtructions having the coincidence. We underline that, by Lemma 3.30, if $g\in\mathrm{GL}_{2}(m)$ is not split liftable modulo $mn$, then all groups containing $g$ are not split liftable modulo $mn$.
@@
In particular, if $E/F$ does not have CM and has a $(p^{k},p^{k+1})$ coincidence, then $E/F$ has non maximal image modulo $p^{2}$, and even modulo $p$ if $p\neq 2,3$.
@@ 4. Large images
For a group $G$, we denote by $\operatorname{\mathrm{D}}(G)$ its commutator subgroup. We know that $D(G)$ is the smallest normal subgroup of $G$ such that $G/\operatorname{\mathrm{D}}(G)$ is abelian and this last is called the abelianization of $G$. We will use several well-known results about derived group of $\mathrm{SL}_{2}(m)$ and $\mathrm{GL}_{2}(m)$, given with detailed proofs in Appendix A.
@@
Now, suppose that $\operatorname{\mathrm{D}}(G)\neq\mathrm{SL}_{2}(m)$. If $3\nmid m$, this does not happens, since in this case $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))=\mathrm{SL}_{2}(m)$ from Proposition A.2. If $3\mid m$, then $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))$ has index $3$ in $\mathrm{SL}_{2}(m)$ from Proposition A.2 and so is $\operatorname{\mathrm{D}}(G)$. It follows that $F(E[m])\cap F^{\mathrm{ab}}$ is an extension of degree $3$ of $F(\zeta_{m})$. ∎
@@
The case $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ happens only for $\gcd(m,12)\neq 1$ by Proposition A.2 and Proposition A.4. Let $k=v_{3}(m)$. In the previous proposition, $m$ is odd and so $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ only if $k>0$. In this case $L:=F(E[3^{k}])\cap F^{\mathrm{ab}}$ is a $(\mathbb{Z}/3\mathbb{Z})$-extension of $F(\zeta_{3^{k}})$ and
@@
For $p\geq 5$, if $\rho_{E,p}(G_{F})$ contains $\mathrm{SL}_{2}(p)$, then $\rho_{E,p^{\infty}}(G_{F})$ contains $\mathrm{SL}_{2}(\mathbb{Z}_{p})$, by [Ser89, IV.3.4.Lemma 3].
@@
Under the hypotheses of the previous theorem, for $(m,n)=(p^{k},p^{k+1})$, we know by the previous theorem that $p=2$ or $3$. Since $\mathrm{SL}_{2}(p^{k})$ contains the matrix $T$, this results was already known by Theorem 3.35 and even more: the assumption $\zeta_{p^{k+1}}\notin F$ is unnecessary and $k=1$.
@@
Let $E/F$ be an elliptic curve with $j$-invariant $j(E)$. We have $F(j(E)^{1/3})\subseteq F(E[3])$. Moreover, if $\mathrm{SL}_{2}(3)\subseteq\rho_{E,3}(G_{F})$, then $F(j(E)^{1/3})/F$ is non trivial.

### v3

@@ Introduction
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{4}$
@@
Let $E/F$ be an elliptic curve with an $(m,n)$-coincidence. Let $p$ be a prime such that $p\mid n$ and $p\nmid m$. Let $\mathfrak{p}$ an ideal above $p$ such that the ramification index of $F/\mathbb{Q}$ at $\mathfrak{p}$ is prime to $\varphi(p^{k})$. Then we are in one of the following cases:
@@ 1. Notations and preliminaries
    For a positive integer $m$, we set
    $\mathrm{GL}_{2}(m):=\mathrm{GL}_{2}(\mathbb{Z}/m\mathbb{Z})\quad\text{and}\quad\mathrm{SL}_{2}(m):=\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z}).$
    For finite extensions $K\subseteq L\subseteq M$ such that $M/L$ is Galois, and a prime ideal $\mathfrak{p}$ of $L$, we denote by $e_{\mathfrak{p}}(M/K)$ the ramification index of $M/K$ at $\mathfrak{p}$.
@@
    For a prime $p$, $\mathbb{Q}(\mu_{p^{\infty}})$ denotes the number field generated by the $\zeta_{p^{k}}$, for $k\geq 1$.
@@
First of all, we prove Proposition 1.2 given in the introduction after some preliminaries. Let $E/F$ be an elliptic curve. If $E/F$ has CM, we suppose that $F$ contains the CM field. From Serre’s open image theorem [Ser72, Section 4.4, Theorem 3’] for the non-CM case, and from [CP22] for the CM case, there exists a finite set $S$ of rational primes such that
@@
Let $m\geq 2$, $p$ be a prime such that $p>q$ for all primes $q\mid m$ and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\mu_{p^{\infty}})$. Let $E/F$ be an elliptic curve such that $F(\zeta_{p^{k}})\subseteq F(E[m])$ with $k>r$. Then, $k=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case either $r=0$ and $k\leq 2$, or $r=k-1$.
@@ 2.2. Ramification and entanglement
The extension $F(\zeta_{p^{k}})/F$ is Galois and so the ramification index above $\mathfrak{p}$ only depends on $\mathfrak{p}$. We have
@@
split multiplicative red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)=0$
additive red. at $\mathfrak{p}$ with $p>3$
@@
(non split) multiplicative red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$
@@
split mult. red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)=0$ Proposition 2.2
add. red. at $\mathfrak{p}$, $p>3$ Proposition 2.3
@@
(non split) mult. red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$ Proposition 2.2
@@
By Lemma 2.5, we obtain the table below, in which we present the necessary condition on the ramification of $F/\mathbb{Q}$ to obtain the ramification index as in previous theorem.
@@
With the notation of Theorem 2.6 and Remark 2.7, if $F(\zeta_{p^{k}})\subseteq F(E[m])$, then we must have $s\mid t$. Therefore, the tables give restrictions on having $F(\zeta_{p^{k}})\subseteq F(E[m])$. For example, if we have this inclusion and $E/F$ has good reduction at $\mathfrak{p}$, then $\varphi(p^{k})$ must divide the ramification index of $F/\mathbb{Q}$ at $\mathfrak{p}$. In the following corollary, we consider the case of $F/\mathbb{Q}$ unramified above $p$.
@@
Again using ramification, we also deduce a result on vertical coincidences, which is the topic of the next section:
Theorem 2.11.
Let $E/F$ be an elliptic curve, $p$ be a prime and $k\geq 2$. If $p^{k-1}\nmid e_{\mathfrak{p}}(F/\mathbb{Q})$ and $E/F$ has good supersingular reduction at $\mathfrak{p}$, then $F(E[p])\neq F(E[p^{k}])$.
Proof.
From [Ser72, Proposition 12], the extension $F(E[p])/F$ is tamely ramified at $p$. Since $p^{k-1}$ does not divide $e_{\mathfrak{p}}(F/\mathbb{Q})$, then Remark 2.7 gives that $F(E(\zeta_{p^{2}})/F$ is wildly ramified at $p$. The result follows as a consequence of the Weil pairing 1.3. ∎
@@ 3.1. Construction of vertical coincidences
$1\to\operatorname{\mathrm{Gal}}(F(E[mn])/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn])/F)\to\operatorname{\mathrm{Gal}}(F(E[m])/F)\to 1$
@@
    $1\to\operatorname{\mathrm{Gal}}(F(E[mn])/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn])/F)\to\operatorname{\mathrm{Gal}}(F(E[m])/F)\to 1$
@@
$1\to(\mathbb{Z}/2\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q})\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[4])/\mathbb{Q})\to 1$
@@
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{r}$
for some $r\leq 4$. Indeed, the Galois group $\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F(E[p^{k}]))$ is isomorphic, for $n=2$, to a subgroup of
@@
$1\to(\mathbb{Z}/p\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F)\to\operatorname{\mathrm{Gal}}(F(E[p^{k}])/F)\to 1$
@@ 3.2. Trivial intersection with the cyclotomic field
If $E/F$ has an $(m,mn)$-coincidence, then we must have the inclusion $F(\zeta_{mn})\subseteq F(E[m])$. But this does not implies in general that $F(\zeta_{mn})=F(\zeta_{m})$, as we will see in Remark 3.27. Even more, unless $m$ is odd and $n=2$, this last never happens if $F=\mathbb{Q}$, and yet some coincidences occurs, like $(2,4)$ and $(2,6)$-coincidence, see [DLR23, Examples 1.2 and 1.3]. As in Remark 3.27, it is due to the non-surjectivity of $\mathrm{SL}_{2}(mn)\cap\rho_{E,mn}(G_{F})\to\mathrm{SL}_{2}(m)\cap\rho_{E,m}(G_{F})$.
@@ 3.3. Index of images
Corollary 3.19.
Let $k\geq 2$ if $p$ is odd and $k\geq 3$ if $p$ is even. Let $E/F$ be an elliptic curve such that the exact sequence
$1\to\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F(E[p^{k}]))\to\operatorname{\mathrm{Gal}}(F(E[p^{k+1}])/F)\to\operatorname{\mathrm{Gal}}(F(E[p^{k}])/F)\to 1$
is split, then $F(E[p^{k-1}])=F(E[p^{k}])$.
Proof.
By Proposition 3.4, there exists $L/F$ linearly disjoint from $F(E[p^{k}])/F$ such that $L(E[p^{k}])=L(E[p^{k+1}])$. Therefore, by Theorem 3.18, we have $L(E[p^{k-1}])=L(E[p^{k}])$. But $L\cap F(E[p^{k}])=F$ gives $L(E[p^{k-1}])\cap F(E[p^{k}])=F(E[p^{k-1}])$, from which we obtain
$[F(E[p^{k}]):F(E[p^{k-1}])]=[L(E[p^{k}]):L(E[p^{k-1}])]=1,$
using [Bou81, A.V.13, Proposition 5]. ∎
Corollary 3.20.
If $p$ is odd or $k\geq 3$, then the sequence
$1\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[p^{k+1}])/\mathbb{Q}(E[p^{k}]))\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[p^{k+1}])/\mathbb{Q})\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[p^{k}])/\mathbb{Q})\to 1$
is not split.
Proof.
If the sequence is split, then the coincidence $\mathbb{Q}(E[p^{k}])=\mathbb{Q}(E[p^{k-1}])$ holds from Corollary 3.19, which is impossible by [DLR23, Theorem 1.4]. ∎
Remark 3.21.
We notice that the assumptions of the theorem are necessary. Indeed, let $E/\mathbb{Q}$ be an elliptic curve satisfying $\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q})\simeq(\mathbb{Z}/2\mathbb{Z})^{6}$. Then, the sequence in Corollary 3.19 is split. However, since $\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q}(E[4]))\leq(\mathbb{Z}/2\mathbb{Z})^{4}$ by (2), we have $(\mathbb{Z}/2\mathbb{Z})^{2}\leq\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[4])/\mathbb{Q})$ and so $\mathbb{Q}(E[2])\neq\mathbb{Q}(E[4])$.
We obtain the following proposition on the adelic index:
Proposition 3.22.
Let $E/F$ be an elliptic curve without CM, with a $(p^{k},p^{k+1})$-coincidence. Then $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ is divisible by $p^{4k}$ if $p$ is odd, or $\mathrm{max}\left\{2^{4},2^{4k-1}\right\}$ if $p$ is even.
Proof.
With the introduced notation, we consider $M=\mathrm{GL}_{2}(\mathbb{Z}_{p})$ and $G=\rho_{E,p^{\infty}}(G_{F})$. The index $i_{k+1}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ by Remark 3.5 and so $i_{k+1}/i_{k}$ divides the global index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$. But having a $(p^{k},p^{k+1})$-coincidence is equivalent to have $G_{k}\simeq G_{k+1}$. Then, the proposition follows from Equation 4 and Corollary 3.17. ∎
Remark 3.23.
In [Zyw22b, Theorem 1.3], Zywina gives a set of possible index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{\mathbb{Q}})]$ which is generic for elliptic curves $E$ defined over $\mathbb{Q}$. In [Zyw24, Theorem 1.2], he gives upper bounds for the index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ for elliptic curves $E/F$, depending on the number field $F$.
@@
    For $E/\mathbb{Q}$ without CM, in [Yvo24, Definition 2.21], we define the $p$-adic depth $s$ as the smallest $k$ such that $\rho_{E,p^{\infty}}(G_{\mathbb{Q}})$ is the full inverse image of $\rho_{E,p^{k}}(G_{\mathbb{Q}})$ in $\mathrm{GL}_{2}(p^{\infty})$. Then, the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ stabilizes at $1$ from $\mathrm{max}\{1,s\}$. In particular, if $G=\mathrm{GL}_{2}(\mathbb{Z}_{p})$, then $\left(\frac{i_{k+1}}{i_{k}}\right)$ is constant equal to $1$. The converse is false: the elliptic curve with LMFDB label 11.a2 has non maximal Galois representation at $5$ and the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ attached to $5$ is constant, equal to $1$. See [Yvo24, Figure 2.1] for the possible $p$-adic depth for any $p$.
@@
Here are some examples over $\mathbb{Q}$, computed from [LMF24], illustrating different possibilities for the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$. Using point (2) of Remark 3.24 and [Yvo24, Table 2.1], the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ is constant equal to $1$ for $p\geq 13$, and it stabilizes at most at rank $5$ if $p=2$, at rank $3$ if $p=3$ and at rank $2$ if $p=5,7,11$.
@@
In the first case, we have seen in Remark 3.26 that $F(\zeta_{p^{k+1}})$ is equal to $F(\zeta_{4})$ if $p=2$ and $F(\zeta_{p})$ otherwise. The examples of vertical coincidence we have for elliptic curves over $\mathbb{Q}$ fit, obviously, in the second case. Indeed, the elliptic curve with LMFDB label 40.a4 has a $(2,4)$-coincidence with
@@ 3.4. Split liftable subgroups
Now we will present groups which are not split liftable, which give us obstructions having the coincidence. We underline that, by Lemma 3.33, if $g\in\mathrm{GL}_{2}(m)$ is not split liftable modulo $mn$, then all groups containing $g$ are not split liftable modulo $mn$.
@@
In particular, if $E/F$ does not have CM and has a $(p^{k},p^{k+1})$-coincidence, then $E/F$ has non maximal image modulo $p^{2}$, and even modulo $p$ if $p\neq 2,3$.
@@ 4. Large images
For a group $G$, we denote by $\operatorname{\mathrm{D}}(G)$ its commutator subgroup. We know that $D(G)$ is the smallest normal subgroup of $G$ such that $G/\operatorname{\mathrm{D}}(G)$ is abelian and this last is called the abelianization of $G$. We will use several well-known results about the derived group of $\mathrm{SL}_{2}(m)$ and $\mathrm{GL}_{2}(m)$, given with detailed proofs in Appendix A.
@@
Now, suppose that $\operatorname{\mathrm{D}}(G)\neq\mathrm{SL}_{2}(m)$. If $3\nmid m$, this does not happen, since in this case $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))=\mathrm{SL}_{2}(m)$ from Proposition A.2. If $3\mid m$, then $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))$ has index $3$ in $\mathrm{SL}_{2}(m)$ from Proposition A.2 and so is $\operatorname{\mathrm{D}}(G)$. It follows that $F(E[m])\cap F^{\mathrm{ab}}$ is an extension of degree $3$ of $F(\zeta_{m})$. ∎
@@
The case $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ occurs only for $\gcd(m,12)\neq 1$ by Proposition A.2 and Proposition A.4. Let $k=v_{3}(m)$. In the previous proposition, $m$ is odd, and so $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ only if $k>0$. In this case $L:=F(E[3^{k}])\cap F^{\mathrm{ab}}$ is a $(\mathbb{Z}/3\mathbb{Z})$-extension of $F(\zeta_{3^{k}})$ and
@@
Let $k=1$ if $p\geq 5$, $k=2$ if $p=3$ and $k=3$ if $p=2$. If $\rho_{E,p^{k}}(G_{F})$ contains $\mathrm{SL}_{2}(p)$, then $\rho_{E,p^{\infty}}(G_{F})$ contains $\mathrm{SL}_{2}(\mathbb{Z}_{p})$, by [Ser89, IV.3.4. Lemma 3].
@@
Under the hypotheses of the previous theorem, for $(m,n)=(p^{k},p^{k+1})$, we know by the previous theorem that $p=2$ or $3$. Since $\mathrm{SL}_{2}(p^{k})$ contains the matrix $T$, this result was already known by Theorem 3.38 and even more: the assumption $\zeta_{p^{k+1}}\notin F$ is unnecessary and $k=1$.
By [Zyw22a, Lemma 1.7], we have $\mathrm{SL}_{2}(\hat{\mathbb{Z}})\cap\rho_{E}(G_{F})=\operatorname{\mathrm{D}}(\rho_{E}(G_{F}))$. In particular, if $\rho_{E}(G_{F})$ contains $\mathrm{SL}_{2}(\hat{\mathbb{Z}})$, then $\operatorname{\mathrm{D}}(\rho_{E}(G_{F}))=\mathrm{SL}_{2}(\hat{\mathbb{Z}})$. In particular, if $8\cdot 9\mid m$ and $\mathrm{SL}_{2}(m)\subseteq\rho_{E,m}(G_{F})$ then $\mathrm{SL}_{2}(\hat{\mathbb{Z}})\subseteq\rho_{E}(G_{F})$ by [KS09, Lemma 1 and the following paragraph], and so $\operatorname{\mathrm{D}}(\rho_{E,m})=\mathrm{SL}_{2}(m)$. Hence we deduce from Theorem 4.1 the following:
Theorem 4.6.
Let $E/F$ be an elliptic curve and $m$ be a positive integer. Suppose that $72\mid m$ and $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(m)$. Then $F(E[m])\cap F^{\mathrm{ab}}=F(\zeta_{m})$. In particular, if $E/F$ has an $(m,n)$-coincidence, then $F(\zeta_{n})\subseteq F(\zeta_{m})$.
@@
Let $E/F$ be an elliptic curve with $j$-invariant $j(E)$. Then we have $F(j(E)^{1/3})\subseteq F(E[3])$. Moreover, if $\mathrm{SL}_{2}(3)\leq\rho_{E,3}(G_{F})$, then the extension $F(j(E)^{1/3})/F$ is non trivial.